# Gamma-ray burst Jet Progenitor: long vs short bursts by angular momentum and mass ratio

### v1

@@ 6.1 Photon-pair sphere/jet temperature profiles
-- $\frac{1}{2}{\mathcal{A}}^{2}a^{2}\Omega r^{2}\sin^{2}\theta\Big{\}}.$
The first term inside the braces corresponds to the gravitational potential energy contribution to the temperature. The second term accounts for the rotational energy’s influence, which decreases the photon-pair plasma temperature due to the repulsive nature of the centrifugal potential $\phi_{c}$ acting against gravitational collapse. For $\theta\not=0$, the photon-pair temperature $T_{\gamma}$ becomes smaller, compared with non-rotating cases $\Omega=0$.
The expression for $T_{\gamma}$ highlights the interplay between the gravitational and centrifugal potentials in determining the thermal energy distribution of the photon-pair plasma during the homologous collapse of the rotating stellar core.
At the end of the homologous collapse $t=t_{\rm final}$, the core’s central density reaches its maximum value $\rho_{c}=\rho_{c}^{\rm max},$ and the sound speed squared is $(v^{c}_{s})^{2}=1/3$, the produced photon-pair plasma forms an axial symmetric jet, described in the cylindrical coordinates. For illustration, we use the parameters, we use parameters $(\gamma,\lambda_{m})=(1.23,8\times 10^{-5})$ (5.22) (the blue case) and ${\mathcal{A}}^{2}a^{2}\Omega=0.05$, indicating a small centrifugal potential $\phi_{c}$.
In Fig. 3, we plot the photon-pair plasma temperature $T_{\gamma}$ for the non-rotating case $\Omega=0$ and rotating cases $\Omega(t)\not=0$. The maximal temperature is located at the centre $r=0$. The vanishing temperature $T_{\gamma}(r,\theta)=0$ determines the photon-pair sphere radius $r_{\gamma}<r_{s}$ or the jet boundary
@@
The photon-pair sphere volume $V_{\gamma}$ is larger than the photon-pair jet volume $V^{j}_{\gamma}$, see Fig. 4. In the rotating case, we plot the jet boundary $r^{j}_{\gamma}(\theta)$ in Fig. 5. As a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than baryon mass ($T_{\gamma}\approx 0.21T\ll m$). It fills the condition for applying the virial theorem (3.2).
Figure 3: Colours online. Using results (6.1) for maximal sound velocity $(v_{s}^{c})^{2}=1/3$, we plot the temperature $T_{\gamma}/m$ in the cylindrical coordinates $y=R\sin\theta$ and $z=R\cos\theta$, where the $\theta$ is the zenith angle. The temperature unit is of typical baryon mass $m\approx 1$ GeV. The parameters are the $Blue$ case parameters (5.22), and the initial conditions are (5.27). The upper panel is a non-rotating case ${\mathcal{A}}^{2}a^{2}\Omega=0$. The lower panel is a rotating case of ${\mathcal{A}}^{2}a^{2}\Omega=0.05$. The temperature vanishes at the photon-pair sphere radius $r_{\gamma}$ or the jet boundary $r^{j}_{\gamma}(\theta)$. Both are smaller than the core outer radius $r_{s}$ (5.22), see Fig. 1.
6.2 Photon-pair sphere/jet density and size profiles
We use photon productions (4.1) and (4.2) by baryon collisions to approximately obtain the photon-pair jet energy and number densities
@@ 6.2 Photon-pair sphere/jet density and size profiles
in terms of photon-pair thermal temperature $T_{\gamma}$. The results show that the photon-pair energy density $\rho_{\gamma}$ is in the range of $(10^{25}\sim 10^{30})~{}{\rm ergs}/{\rm cm}^{3}$, and the number density $n_{\gamma}$ is in the range of $(10^{37}\sim 10^{30})/{\rm cm}^{3}$.
@@
and the photon-pair jet boundary
@@
They are expressed as a function of initial core centre density $\rho^{\rm in}_{c}$ when a homologous collapse initiates. The photon-pair sphere size $R_{\gamma}\sim(10^{7}\sim 10^{10})~{}{\rm cm}$ and the photon-pair cylindrical jet width $R^{j}_{\gamma}(\pi/2)\sim(10^{6}\sim 10^{9})~{}{\rm cm}$.
Both sizes $R_{\gamma}$ and $R^{j}_{\gamma}$ are much larger than the photon mean-free path $\lambda_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$, namely $\sigma_{\gamma}n_{\gamma}R_{\gamma}\gg 1$ or $\sigma_{\gamma}n_{\gamma}R^{j}_{\gamma}\gg 1$. The photon-pair sphere/jet is deeply opaque. On the other hand, we find in Fig. 3 that the photon-pair sphere/jet temperature $T_{\gamma}$ can be well above the critical energy threshold $1.022m_{e}=$ 1.02 MeV of electron-positron ($e^{+}e^{-}$) pair production. The photon-pair sphere/jet energy density $\rho_{\gamma}$ can be well above the critical energy density $\rho_{e}=m_{e}^{4}=5.93\times 10^{-11}\rho_{0}$ of electron-positron pairs. This means the photons are quickly thermalised to form an electron-position-photon plasma, as briefly discussed in Sec. 3. The photon-pair sphere/jet undergoes ultra-relativistic hydrodynamic expansion, which has been well-studied in the literature, and will not be studied in this article.
We compare the photon-pair sphere and jet boundaries in Fig. 4 and define the cylindrical jet angle $\theta_{\rm jet}$ as the ratio
$\theta_{\rm jet}\approx\frac{R^{j}_{\gamma}(\theta=\pi/2)}{R^{j}_{\gamma}(\theta=0)},\quad R^{j}_{\gamma}(\theta=0)=R_{\gamma},$ (6.6)
that characterizes the axial symmetric jet along the $\hat{z}$ direction.
Figure 4: We compare the photon-pair sphere and jet boundaries of Fig. 3. The photon-pair sphere radius $r_{\gamma}\approx 30$ and $R_{\gamma}=a_{\rm min}r_{\gamma}\sim(10^{7}-10^{10})$ cm. The photon-pair jet boundary $r^{j}_{\gamma}(\theta)$ is given in Fig. 5.
Figure 5: The photon-pair jet boundary $r^{j}_{\gamma}(\theta)$ shows that (i) $\theta\rightarrow 0$ and $r^{j}_{\gamma}(\theta)$ approaches to the radius $r_{\gamma}\approx 30$ of the photon-pair sphere; (ii) $\theta\rightarrow\pi/2$ and $r_{\gamma}(\theta)$ approaches to $r^{j}_{\gamma}(\pi/2)\approx 1$ of the photon-pair jet. The cylindrical jet width $R^{j}_{\gamma}=a_{\rm min}r^{j}_{\gamma}(\pi/2)\sim(10^{6}-10^{8})$ cm. The jet angle (6.6) $\theta_{\rm jet}\approx 1/30\approx 1^{\circ}$.
6.3 Photon-pair sphere/jet energy and spectrum peak
The total relativistic particle energy $E_{\gamma}$ and number $N_{\gamma}$ of the photon-pair sphere/jet can be obtained by numerically integrating the energy and number densities (6.3) over the photon-pair jet volume
@@ 6.3 Photon-pair sphere/jet energy and spectrum peak
over photon-pair jet volume $V^{j}_{\gamma}$. From Eq. (6.3), we obtain the total energy and photon number of photon-pair jet systems
@@
In addition, we define the mean temperature of the photon-pair sphere/jet as averaged value of $T_{\gamma}$ (6.1) over photon-pair jet volume $V^{j}_{\gamma}$,
@@
which represents the characteristic temperature of the photon-pair sphere/jet. The photon-pair spectrum has a maximum at the peak energy $E_{p}\sim\langle T_{\gamma}\rangle$.
In Eqs. (6.10) and (6.12), we express total photon-pair energy $E_{\gamma}$ and spectral peak $E_{p}$ in terms of the initial core centre density $\rho^{\rm in}_{c}=\rho_{c}(t_{\rm initial})$ at the time $t_{\rm initial}$ when the core initiates homologous collapse. If the core mass $M$ and baryon number $N$ are fixed and conserved, the smaller initial core centre density $\rho^{\rm in}_{c}$ is, the more extended core mass density profile $\rho(r,t_{\rm initial})$ is, i.e., the larger $r_{s}$ is. As a consequence, more gravitational energy can be gained in collapse and converted to the photon-pair sphere/jet energy $E_{\gamma}$ and number $N_{\gamma}$. In contrast, these quantities are also functions of the material binding energy parameter $\lambda$, the averaged thermal index $\gamma$ and rotation parameter $J/M\propto{\mathcal{A}}^{2}a^{2}\Omega$. The rotation represents the repulsive reactions against gravitational collapse and energy gain.
@@
where the variations are due to the ranges of the parameters $\lambda$, $\gamma$ and $J/M\propto{\mathcal{A}}^{2}a^{2}\Omega$. These quantities are conserved in the subsequent processes of the ultra-relativistic hydrodynamic expansion and internal/external collisions up to the transparent point when pairs annihilate and produce photons stream to detectors.

### v2

@@ 6.2 Collimated fireball temperature of photons and $e^{+}e^{-}$ pairs
-- $\frac{1}{2}a^{2}\Omega^{2}_{c}r^{2}\sin^{2}\theta\Big{\}},$
at the end of homologous collapses, when $t=t_{\rm final}=\tau_{\rm min}$, minimal value $a(\tau_{\rm min})=a_{\rm min}$ (5.22) in Fig. 2, maximal values $[v^{c}_{s}(\tau_{\rm min})]^{2}=(v^{c}_{s})^{2}=1/3$, $\Omega_{c}(\tau_{\rm min})=\Omega_{c}$ and $a^{2}\Omega_{c}^{2}\equiv a^{2}(\tau_{\rm min})\Omega^{2}_{c}(\tau_{\rm min})$. The first term inside the braces corresponds to the gravitational potential energy contribution to the temperature. The second term accounts for the rotational energy’s influence, which decreases the photon-pair temperature due to the repulsive nature of the centrifugal potential $\phi_{c}$ acting against gravitational collapse. For $\theta\not=0$, the photon-pair temperature $T_{\gamma}$ becomes smaller, compared with non-rotating cases $\Omega_{c}=0$. Photon-pair fireballs are asymmetrically collimated along the rotation axes.
The expression for $T_{\gamma}$ highlights the interplay between the gravitational and centrifugal potentials in determining the thermal energy distribution of the photon-pair plasma during the homologous collapse of the rotating stellar core. We describe the produced photon-pair plasma forms axial symmetric collimation in the cylindrical coordinates. For illustration, we use the parameters, we use parameters $(\gamma,\lambda_{m})=(1.23,8\times 10^{-5})$, the blue case in the list (5.20), and $a^{2}\Omega^{2}_{c}=0.05$ for a slowly rotating core.
In Fig. 4, we plot the photon-pair temperature $T_{\gamma}$ for the non-rotating case $\Omega_{c}=0$ and rotating cases $\Omega_{c}\not=0$. The maximal temperature is located at the centre $r=0$. The temperature $T_{\gamma}(r,\theta)=2m_{e}$ determines the photon-pair sphere fireball radius $r_{\gamma}<r_{s}$ or the collimated fireball boundary
@@
The photon-pair sphere volume $V_{\gamma}$ is larger than the photon-pair collimated volume $V^{j}_{\gamma}$, see Fig. 5. In rotating cases, we plot the collimated boundary $r^{j}_{\gamma}(\theta)$ in Fig. 6. As a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than baryon mass ($T_{\gamma}\approx 0.21T\ll m$). It fills the condition for applying the virial theorem (3.3).
Figure 5: Following the colourmap of Fig. 4 for spherical and collimated fireballs, we compare the photon-pair fireball radius $r_{\gamma}$ and collimated fireball boundary $r^{j}_{\gamma}(\theta)$ (6.6), see Fig. 4, with the boundary $r_{s}\approx 36$ of the baryon core density $f^{3}(r)$ (5.12), see Fig. 1. The photon-pair fireball radius $r_{\gamma}\lesssim 30$ and $R_{\gamma}=a_{\rm min}r_{\gamma}\sim(10^{7}-10^{10})$ cm. The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ is given in Fig. 6. Note that the green zone and boundary indicate a baryon core, the blue (orange) zone and boundary indicate a spherical (collimated) fireball of photons and $e^{+}e^{-}$ pairs, and the boundary defines at the threshold $T_{\gamma}(r_{\gamma})=2m_{e}$ ($T_{\gamma}[r^{j}_{\gamma}(\theta)]=2m_{e}$) respectively. The baryon temperature zone and boundary follow the photon-pair temperature zone and boundary for $T_{\gamma}=0.21T$. These spatial distributions of fireballs and baryon core at the end of homologous collapses show $r_{s}>r_{\gamma}\geq r^{j}_{\gamma}(\theta)$. The collimated fireballs (orange zone) rotate along $\hat{z}$ axes with nontrivial angular momentum $J_{\gamma}$ (7.1). They act as initial configurations for ultra-relativistic hydrodynamic evolution (early launching and later propagation): Spherical expansion or axial-symmetric jet evolution outwards, proceeding internal and external shocks to produce GRBs.
Figure 6: The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ shows that (i) $\theta\rightarrow 0$ and $r^{j}_{\gamma}(\theta)$ approaches to the radius $r_{\gamma}\lesssim 30$ of the photon-pair fireball; (ii) $\theta\rightarrow\pi/2$ and $r_{\gamma}(\theta)$ approaches to $r^{j}_{\gamma}(\pi/2)\approx 1$ of the photon-pair collimated fireball (the cylindrical collimation) width $R^{j}_{\gamma}=a_{\rm min}r^{j}_{\gamma}(\pi/2)\sim(10^{6}-10^{8})$ cm, which is the maximal lateral size. The collimated fireball angle (7.2) $\theta_{\rm coll}\approx 1.8/30\approx 3.4^{\circ}$.
6.3 Collimated fireball energy density and size profiles
We use photon productions (6.2) and (6.3) by baryon collisions to approximately obtain the collimated fireball energy and number densities of photons and pairs
@@ 6.3 Collimated fireball energy density and size profiles
in terms of collimated fireball temperature $T_{\gamma}(r,\theta)$. The results show that the fireball energy density $\rho_{\gamma}$ is in the range of $(10^{25}\sim 10^{30})~{}{\rm ergs}/{\rm cm}^{3}$, and number density $n_{\gamma}$ is in the range of $(10^{37}\sim 10^{30})/{\rm cm}^{3}$.
@@
and the photon-pair collimation boundary
@@
They are expressed as a function of initial core centre density $\rho^{\rm in}_{c}$ when a homologous collapse initiates. The photon-pair sphere fireball size $R_{\gamma}\sim(10^{7}\sim 10^{10})~{}{\rm cm}$ and the photon-pair cylindrical collimation width $R^{j}_{\gamma}(\pi/2)\sim(10^{6}\sim 10^{9})~{}{\rm cm}$.
Both sizes $R_{\gamma}$ and $R^{j}_{\gamma}$ are much larger than the photon mean-free path $\lambda_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$, namely $\sigma_{\gamma}n_{\gamma}R_{\gamma}\gg 1$ or $\sigma_{\gamma}n_{\gamma}R^{j}_{\gamma}\gg 1$. The photon-pair spherical and collimated fireballs are deeply opaque. On the other hand, we find in Fig. 4 that the photon-pair spherical and collimated fireballs’ temperature $T_{\gamma}$ can be well above the critical energy threshold $1.022m_{e}=$ 1.02 MeV of electron-positron ($e^{+}e^{-}$) pair production. The photon-pair spherical/collimated fireball energy densities $\rho_{\gamma}$ can be well above the critical energy density $\rho_{e}=m_{e}^{4}=5.93\times 10^{-11}\rho_{0}$ of electron-positron pairs. This means the photons are quickly thermalised to form an electron-position-photon plasma, as briefly discussed in Sec. 3.
Up to now, we have studied in this simplified model the processes that the gravitational binding energy is extracted and converted to a collimated and opaque fireball energy at the end of gravitational stellar collapses. The effect of an external magnetic field in such a process needs further investigation. These studies of extracting gravitational binding energy in stellar collapses contrast with the processes [22, 23, 24] of extracting rotational energy from spinning black holes.
6.4 Collimated fireball energy and spectrum peak
The total relativistic particle energy $E_{\gamma}$ and number $N_{\gamma}$ of the photon-pair collimated fireball can be obtained by numerically integrating the energy and number densities (6.3) over the photon-pair collimated volume
@@ 6.4 Collimated fireball energy and spectrum peak
over photon-pair collimated volume $V^{j}_{\gamma}$. From Eq. (6.3), we obtain the total energy and photon number of photon-pair collimated systems
@@
In addition, we define the mean temperature of the photon-pair sphere/collimation as averaged value of $T_{\gamma}$ (6.5) over photon-pair collimated volume $V^{j}_{\gamma}$,
@@
which represents the characteristic temperature of the photon-pair sphere/collimation. The photon-pair spectrum has a maximum at the peak energy $E_{p}\sim\langle T_{\gamma}\rangle$.
In Eqs. (6.13) and (6.15), we express total photon-pair energy $E_{\gamma}$ and spectral peak $E_{p}$ in terms of the initial core centre density $\rho^{\rm in}_{c}=\rho_{c}(t_{\rm initial})$ at the time $t_{\rm initial}$ when the core initiates homologous collapse. The final core centre density is about 10 times the nuclear density $\rho^{\rm max}_{c}\approx 10\rho_{0}$. If the core mass $M$ and baryon number $N$ are fixed and conserved, the smaller initial core centre density $\rho^{\rm in}_{c}$ is, the more extended core mass density profile $\rho(r,t_{\rm initial})$ is, i.e., the larger $r_{s}$ is. As a consequence, more gravitational energy can be gained in collapse and converted to the photon-pair sphere/collimation energy $E_{\gamma}$ and number $N_{\gamma}$. In contrast, these quantities are also functions of the material binding energy parameter $\lambda$, the averaged thermal index $\gamma$ and rotation parameter ${\mathcal{A}}^{2}a^{2}\Omega_{c}^{2}\propto J/M$. The rotation represents the repulsive reactions against gravitational collapse and energy gain.
@@
where the variations are due to the ranges of the parameters $\lambda$, $\gamma$, and small values $a^{2}\Omega_{c}^{2}\propto\Omega_{c}(J/M)$. These results are consistent with the total energy and entropy budgets required to explain the cosmological origin of the GRB phenomena observed. We recall that in the Newtonian approximation, the variation of gravitational energy $GM^{2}/(2R)$ is $M/4$ for a core-collapse from infinity $R=\infty$ to $R=2GM$.
The realistic situations of baryon collisional relaxation and photon production in gravitational collapse and hydrodynamic processes are much more complicated. Among others, we mention a few questions about internal perturbations and shock waves formed due to the baryon matter EoS (3.4) and heat function variations, phase transition and material binding energy change as nuclear matter density variations, fluid plasma stability in external magnetic fields and general relativistic effects. We expect the reduction of collimated fireball energy and number densities, the softened energy spectra, and the prolonged collapsing time. These questions will be the subject of future studies.
7 Collimated fireball angular momentum and relativistic jet
7.1 Angular momentum conservation making an ultra-relativistic jet
The collimated photon-pair fireball is neutral and free from the classical forces of large-scale electromagnetic fields since there are no macroscopic distributions of electric charges and currents. In addition, the collimated fireball carries a nontrivial angular momentum $J_{\gamma}$ along the $\hat{z}$ direction in the rest frame,
$J_{\gamma}/M=M^{-1}\int d^{3}R\rho_{\gamma}j,\quad J_{\gamma}/M\ll J/M,\quad J_{\gamma}\propto J$ (7.1)
analogously to the total core angular momentum $J$ calculations (5.7). The $jdm_{\gamma}=j\rho_{\gamma}dV$ is an elementary angular momentum of the photon-pair fluid element of energy $dm_{\gamma}=\rho_{\gamma}dV$ inside a collimated fireball. Since the fireball thermal temperature $T_{\gamma}$ is much smaller than the baryon mass scale $m$, $T_{\gamma}\ll m$ shown in Fig. 4, the collimated fireball energy density $\rho_{\gamma}\propto T_{\gamma}^{4}$ (6.3) is much smaller than the baryon core density $\rho_{\gamma}\ll\rho$, therefore $J_{\gamma}\ll J$. Equations (5.7) and (7.1) show the collimated fireball angular momentum $J_{\gamma}$ is proportional to the angular momentum $J$ of massive rotating systems in gravitational collapses. The total angular momentum conservation in gravitational collapsing processes implies that collimated fireballs acquire angular momentum $J_{\gamma}$ from rotating systems in collapses.
The hydrodynamical outflows of the collimated fireballs due to the huge photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The hydrodynamical outflows of the collimated fireballs due to the extremely high photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The collimated fireball outflow carries away the energy $E_{\gamma}$ and angular momentum $J_{\gamma}$ from collapsed systems, engulfing the outer material and forming an ultra-relativistic ejecta interacting with stellar matter, as described in the standard GRB scenario. The novelty is that the collimated outflows and ejecta launch ultra-relativistic jets of nontrivial angular momenta $J_{\gamma}$ of collimated fireballs. In later propagation, ultra-relativistic jets maintain collimated structure due to the conservation of their nontrivial angular momentum $J_{\gamma}\not=0$, analogously to a gyroscope effect. Numerical/simulation-based studies are necessary for verifying such dynamics. Analytical studies of ultra-relativistic jet launching and propagation and its stability in magnetic fields will be topics for separate articles. Notwithstanding, we present in the next section some qualitative discussions on the ultra-relativistic jet formation, early launching and later propagation.
In contrast, hydromagnetic flows from accretion discs transport away the rotational energy and angular momentum along large-scale toroidal magnetic fields, making a relativistic jet [25].
7.2 Collimated fireball angle $\theta_{\rm coll}$ relates to jet angle
As depicted in Figs. 4 and 5, the collimated fireball of photon-pair plasma energy density $\rho_{\gamma}$ shows denser and wider at the base $(\theta\rightarrow\pi)$, while less dense and narrower at the top $(\theta\rightarrow 0)$. It is the consequence of gravitational binding potential $\phi$ and centrifugal potential $\phi_{c}$ in the temperature profile $T_{\gamma}$ (6.5) at the end of homologous collapses. The opaque collimated fireball pressure (density) $p_{\gamma}\sim\rho_{\gamma}\propto T_{\gamma}^{4}$ is no longer isotropic and its maximal value along $\hat{z}$-axis, decreasing from $R=0$ to $R=R_{\gamma}$. In the transverse directions perpendicular to $\hat{z}$, its value decreases to the collimated fireball boundary (orange thick line in Fig. 5). Namely, most collimated fireball energy power concentrates at the centre along $\hat{z}$.
On the contrary, the collimated fireball’s angular momentum $j\rho_{\gamma}dV$ vanishes along $\hat{z}$ direction ($\theta=0$) from $R=0$ to $R_{\gamma}$, and increases in the transverse directions ($\theta=\pi/2$). Namely, most fireball angular momentum is carried near the collimated fireball boundary. An outward thermal pressure in the fireball centre along the $\hat{z}$ direction and the conservation of angular momentum most near the collimated fireball boundary possibly leads to an ultra-relativistic jet.
In the phase of early jet launching, we argue that such a collimated fireball will self-similarly evolve to the normal jet profile of energy density denser and wider at the top $(\theta\rightarrow 0)$, less dense and narrower at the base $(\theta\rightarrow\pi)$. The reason is that the highly dense and energetic part is deeply opaque and has a large thermal pressure $p_{\gamma}$ along the $\hat{z}$ direction, ultra-relativistically moving outward. As a result, denser and more energetic parts pile up on the front, moving at almost light speed, while dilute and less energetic parts leave behind as a narrow tail. It is also the case for the collimated fireball outflow engulfing a small amount of material. This situation is analogous to the spherical case of ultra-relativistically hydrodynamic expansion: a spherical fireball evolves to a self-similar (homologous) spherical slab at the front moving at almost light speed, and a tail of dilute number and energy densities left behind, which are studied both analytically and numerically [18, 19].
We must mention the condition of a collimated fireball forming an outward propagating jet. Figure 5 shows collimated fireballs form inside external baryon shells (green layers). Whether or not they can erupt out of external baryon shells requires the ratio $E_{\gamma}/M_{\rm ej}\gg 1$, see the references and discussions for Eq. (8.4), where $E_{\gamma}$ is the total energy of collimated fireballs and $M_{\rm ej}$ is the ejecta mass contributing from the external baryon shell. The $M_{\rm ej}$ is much smaller than the total baryon core mass $M$, i.e., $M_{\rm ej}\ll M$, because Figure 1 shows the baryon external shell mass density is very small for $r>10$ comparing the core centre density. While collimated fireball energies are about $E_{\gamma}/M\sim 10^{-2}$ (6.16). This implies the possibility of collimated fireballs erupting from baryon cores and forming ultra-relativistic ejecta with nontrivial angular momenta.
In the phase of later jet propagation, we argue that such a formed ultra-relativistic jet maintains its structure and dynamics due to the conservation of its jet angular momentum $J_{\gamma}$ along the jet propagating direction. The jet angle $\theta_{\rm jet}\approx dR_{\perp}/dR$ can be defined with its radius $R$ and lateral size $R_{\perp}$. However, one has to consider the ultra-relativistic jet interaction with ambient stellar media and its stability in external large-scale magnetic fields. These issues necessarily require numerical studies.
Nevertheless, we compare the photon-pair sphere and collimated boundaries in Fig. 5 and define the cylindrical collimated fireball angle $\theta_{\rm coll}$ as the ratio
$\theta_{\rm coll}\approx\frac{R^{j}_{\gamma}(\theta=\pi/2)}{R^{j}_{\gamma}(\theta=0)},\quad R^{j}_{\gamma}(\theta=0)=R_{\gamma},$ (7.2)
where $R^{j}_{\gamma}(\theta=\pi/2)$ is the maximal lateral size of collimated fireballs and $R^{j}_{\gamma}(\theta=0)$ is the corresponding radius of the spherical fireball. Such an angle $\theta_{\rm coll}$ characterizes the axial symmetric collimation of photon-pair fireball along the $\hat{z}$ direction before its ultra-relativistic hydrodynamic outflow.
The collimated fireball angle $\theta_{\rm coll}$ (7.2) differs from the ultra-relativistic outflow jet angle $\theta_{\rm jet}$ and the observational jet angle $\theta^{*}_{\rm jet}$ defined through the maximal lateral size of the equal arrival time surface in the structure and dynamics of GRB jets [43]. The collimated structure of the ultra-relativistic outflow should be maintained in jet launching and propagation due to the conservation of the angular momentum $J_{\gamma}$ along the jet direction. Therefore, we speculate that the GRB jet angles $\theta_{\rm jet}$ and $\theta^{*}_{\rm jet}$ should proportionally relate (or positively correlate) to the collimated fireball angle $\theta_{\rm coll}$,
$\theta_{\rm coll}\sim\theta_{\rm jet}\sim\theta^{*}_{\rm jet}.$ (7.3)
The smallness of the collimated fireball angle $\theta_{\rm coll}$ implies the smallness of the GRB jet angles $\theta_{\rm jet}$ and $\theta^{*}_{\rm jet}$.
8 Collimated fireball intrinsic correlations
In this simplified model, we obtain the configurations of photon-pair collimated fireballs determined by
$(J/M,\rho^{\rm in}_{c}/\rho_{0},\gamma,\lambda),$ (8.1)